OpenStudy (anonymous):

how do you find the lenghth of the hypotenuse in a 45 degree- 45 degree- 90 degree triangle??

OpenStudy (amistre64):

1^2 + 1^2 = c^2 1+1 = c^2 2 = c^2 sqrt(2) = c

OpenStudy (amistre64):

any rt triangle can be used with the pythag thrm

OpenStudy (anonymous):

 using this picture